3) Prove by induction that for all natural numbers n:


(d) 1/(1)(3) + 1/(3)(5) + · · · +1/(2n − 1)(2n + 1) = n/(2n + 1)




1
Expert's answer
2021-05-31T06:36:18-0400

Step 1: Given data


P(n) : "\dfrac{1}{1.3}+\dfrac{1}{3.5}+−−−−−+\dfrac{1}{(2n−1)(2n+1)}=\dfrac{n}{(2n+1)}"


Step 2: To prove, "P(1)" is true


 "LHS=\dfrac{1}{1.3}=\dfrac{1}{3}"


"RHS=\dfrac{1}{2×1+1}=\dfrac{1}{3}"


Thus, "P(1)" is true.


Step 3: Assume that the statement is true for "n = k" , where k is some positive integer.


"\dfrac{1}{1.3}+\dfrac{1}{3.5}+−−−−−+\dfrac{1}{(2k−1)(2k+1)}=\dfrac{k}{(2k+1)}"

 

Step 4:  For "n=k+1"  


"LHS= \dfrac{1}{1.3}+\dfrac{1}{3.5}+−−−−+\dfrac{1}{(2k−1)(2k+1)}+\dfrac{1}{(2(k+1)−1)(2(k+1)+1)}"


"=\dfrac{k}{(2k+1)}+\dfrac{1}{(2(k+1)−1)(2(k+1)+1)}"     


 "=\dfrac{k}{(2k+1)}+\dfrac{1}{(2k+1)(2k+3)}"


"=\dfrac{2k^2+3k+1}{(2k+1)(2k+3)}"


"=\dfrac{(2k+1)(k+1)}{(2k+1)(2k+3)}"


"LHS=\dfrac{(k+1)}{(2k+3)}"

  

Step 5:

"RHS=\dfrac{(k+1)}{(2(k+1) + 1)}"


"=\dfrac{k+1} { 2k + 3}"


⇒LHS=RHS